# DLoRA-TrOCR: Mixed Text Mode Optical Character Recognition Based On Transformer

### v2

@@ Abstract
With the continuous development of OCR technology and the expansion of application fields, text recognition in complex scenes has become a key challenge. Factors such as multiple fonts, mixed scenes and complex layouts seriously affect the recognition accuracy of traditional OCR models. Although OCR models based on deep learning have performed well in specific fields or similar datasets in recent years, the generalization ability and robustness of the model are still a big challenge when facing complex environments with multiple scenes. Furthermore, training an OCR model from scratch or fine-tuning all parameters is very demanding on computing resources and inference time, which limits the flexibility of its application. This study focuses on a fundamental aspect of mixed text recognition in response to the challenges mentioned above, which involves effectively fine-tuning the pre-trained basic OCR model to demonstrate exceptional performance across various downstream tasks. To this end, we propose a parameter-efficient mixed text recognition method based on pre-trained OCR Transformer, namely DLoRA-TrOCR. This method embeds DoRA into the image encoder and LoRA into the internal structure of the text decoder, enabling efficient parameter fine-tuning for downstream tasks. Experimental results show that compared to similar parameter adjustment methods, our model DLoRA-TrOCR has the smallest number of parameters and performs better. It can achieve state-of-the-art performance on complex scene datasets involving simultaneous recognition of mixed handwritten, printed and street view texts.
@@ 1 Introduction
Optical Character Recognition (OCR) is the technology of classifying optical patterns corresponding to letters, digits, or other characters in digital images[3]. It typically consists of two tasks: text detection and text recognition. Text detection involves detecting text regions from an image, while text recognition extracts and recognizes the text within those regions.
@@ 2 Related Work
OCR models usually adopt a pipeline structure composed of multiple independent modules, such as CNN (Convolutional Neural Network)[8] for feature extraction, LSTM (Long Short Term Memory)[36] for sequence modeling, etc. This compartmentalized design has many flaws, making it difficult for the model to fully utilize global information and long-range dependencies. In recent years, the Transformer model has achieved excellent performance in sequence modeling tasks due to its powerful self-attention mechanism and end-to-end architecture design[4].It mainly uses an encoder-decoder architecture to achieve from image to text in OCR tasks.
@@ 3.1 Transformer-based OCR
The image encoder we adopt is based on the ViT model, which receives the input image $x_{\mathrm{img}}\in\mathbb{R}^{3\times H_{0}\times W_{0}}$ and resizes it to a fixed size $(H,W)$. Then, the input image is divided into $N$ image blocks of fixed size (P, P), tiled into vectors, and linearly projected to D-dimensional vectors. To preserve the position information of each pixel in the encoded image, absolute position embeddings are added to the embedding vector of each image patch. Next, the sequence of vectors is fed into a series of identical Transformer encoder layers. Each Transformer encoder layer contains a self-attention layer and a feed-forward layer. After each layer, a normalization layer and residual connections are applied. The self-attention mechanism calculates the similarity between each query and all keys, and uses the similarity as a weight to weight all values, and finally obtains the output. The feedforward layer usually consists of two fully connected layers and uses an activation function (such as ReLU and Swish[14]) for nonlinearization.
$\operatorname{Self-Att}(Q,K,V)$ $=V\operatorname{softmax}\big{(}\frac{K^{\top}Q}{\sqrt{D_{k}}}\big{)},$
$\operatorname{MultiHead}(Q,K,V)$ $=W_{o}[\operatorname{head}_{1};\cdots;\operatorname{head}_{h}],$
$\operatorname{head}_{i}$ $=\operatorname{Self-Att}(Q_{i},K_{i},V_{i}),$
$\forall i\in\{1,\cdots,h\},\quad Q_{i}$ $=W_{q}^{i}Q,K_{i}=W_{k}^{i}K,V_{i}=W_{v}^{i}V.$
@@
$y_{tw}$ $=Proj(Emb(w_{t}))$
$\text{Softmax}(y_{tw})$ $=\frac{e^{y_{tw}}}{\sum_{v=1}^{V}e^{y_{tv}}}\quad\text{for }w=1,2,...,V$
Here $V$ denotes the vocabulary size, $Emb(w_{t})$ denotes the embedding vector of the word $w_{t}$, and $Proj(\cdot)$ is the linear transformation that maps the embedding vector to the hidden state space.
@@ 3.2 Parameter-efficient Fine-tuning
$W=W_{0}+BA$
where $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are two small matrices, called LoRA weights, which are parameters that need to be learned during the fine-tuning process. $r$ is a hyperparameter that controls the rank of LoRA weight. When $r\ll\min(m,n)$, the number of parameters required by the LoRA weight will be much smaller than the original model weight matrix. By adding this low-rank update, LoRA enables efficient fine-tuning of pre-trained models. During the inference phase, the fine-tuned LoRA weights are combined with the original model weights to obtain the final inference weights. Since only two small matrices are stored, LoRA greatly reduces the additional parameter storage space required for fine-tuning. In our text decoder, we apply LoRA to both the query, key and value projection matrices of the self-attention module of the decoder and the encoder-decoder self-attention module, as well as the linear layer of the output. This parameter-efficient fine-tuning method allows us to significantly reduce the number of parameters required to fine-tune the text decoder while maintaining performance. Specifically, the low-rank matrix of LoRA is used to reconstruct the multi-head attention mechanism in the Transformer block, and its incremental update can be expressed as
$\operatorname{head}_{i}=\operatorname{self-att}({Q}_{i},{K}_{i},{V}_{i})$
$Q_{i}=(W_{q,i}+B_{q,i}A_{q,i})\mathbf{x}$
$K_{i}=(W_{k,i}+B_{k,i}A_{k,i})\mathbf{x}$
$V_{i}=(W_{v,i}+B_{v,i}A_{v,i})\mathbf{x}.$
@@
$O=(W_{o}+B_{o}A_{o})\operatorname{Concat}(\operatorname{head}_{1};\cdots;\operatorname{head}_{h})\$
@@
$\Delta W=\mathbf{u}\frac{V+\Delta V}{\|V+\Delta V\|_{c}}=\|W_{0}\|_{c}\frac{W_{0}+BA}{\|W_{0}+BA\|_{c}}$
@@
$Q_{i}$ $=(W_{q,i}+\|W_{q,i}\|_{c}\frac{W_{q,i}+B_{q,i}A_{q,i}}{\|W_{q,i}+B_{q,i}A_{q,i}\|_{c}})\mathbf{x}$
$K_{i}$ $=(W_{k,i}+\|W_{k,i}\|_{c}\frac{W_{k,i}+B_{k,i}A_{k,i}}{\|W_{k,i}+B_{k,i}A_{k,i}\|_{c}})\mathbf{x}$
$V_{i}$ $=(W_{v,i}+\|W_{v,i}\|_{c}\frac{W_{v,i}+B_{v,i}A_{v,i}}{\|W_{v,i}+B_{v,i}A_{v,i}\|_{c}})\mathbf{x}.$
@@
$\nabla_{A}\mathcal{L}=B^{\top}\frac{\partial\mathcal{L}}{\partial V^{\prime}},\nabla_{B}\mathcal{L}=\frac{\partial\mathcal{L}}{\partial V^{\prime}}A^{\top}$
where
$\frac{\partial{\mathcal{L}}}{\partial{V^{\prime}}}=\frac{\mathbf{u}}{\|V^{\prime}\|_{c}}\left(I-\frac{V^{\prime}V^{\prime\top}}{\|V\|_{c}^{2}}\right)\frac{\partial{\mathcal{L}}}{\partial{W}}$
@@
Table 1: Optimization of computational resources and model trainable parameters using LoRA, DoRA and DLoRA methods, which load the pre-trained weights of the model using fp16.
@@
  • 1. Full LoRA or DoRA: the encoder and decoder both adopt LoRA or DoRA method;
  • 2. Hybrid configuration: the encoder uses DoRA method, while the decoder adopts LoRA method;
  • 3. Full parameter fine-tuning: no PEFT method, all parameters participate in the training.
@@ 4.1 Experiment Setup
Precision=Correct matchesThe number of detected $\mathrm{Precision}=\frac{\text{Correct matches}}{\text{The numberof detected words}}$
Recall=Correct matchesThe number of ground truth $\mathrm{Recall}=\frac{\text{Correct matches}}{\text{The number ofground truth words}}$
@@ 4.2 Experiment Result
Table 3: Different models were simultaneously evaluated in IAM,SROIE, STR benchmark and Mixed Dataset average CER and average F1 score.

### v3

@@ Abstract
With the continuous development of Optical Character Recognition (OCR) and the expansion of application fields, text recognition in complex scenes has become a key challenge. Factors such as multiple fonts, mixed scenes and complex layouts seriously affect the recognition accuracy of traditional OCR models. Although OCR models based on deep learning have performed well in specific fields or similar datasets in recent years, the generalization ability and robustness of the model are still a big challenge when facing complex environments with multiple scenes. Furthermore, training an OCR model from scratch or fine-tuning all parameters is very demanding on computing resources and inference time, which limits the flexibility of its application. This study focuses on a fundamental aspect of mixed text recognition in response to the challenges mentioned above, which involves effectively fine-tuning the pre-trained basic OCR model to demonstrate exceptional performance across various downstream tasks. To this end, we propose a parameter-efficient mixed text recognition method based on pre-trained OCR Transformer, namely DLoRA-TrOCR. This method embeds DoRA into the image encoder and LoRA into the internal structure of the text decoder, enabling efficient parameter fine-tuning for downstream tasks. Experiments show that compared to similar parameter adjustment methods, our model DLoRA-TrOCR has the smallest number of parameters and performs better. It can achieve state-of-the-art performance on complex scene datasets involving simultaneous recognition of mixed handwritten, printed and street view texts.
@@ 1 Introduction
OCR is the technology of classifying optical patterns corresponding to letters, digits, or other characters in digital images[3]. It typically consists of two tasks: text detection and text recognition. Text detection involves detecting text regions from an image, while text recognition extracts and recognizes the text within those regions.
@@ 2 Related Work
OCR models usually adopt a pipeline structure composed of multiple independent modules, such as Convolutional Neural Network(CNN)[9] for feature extraction, Long Short Term Memory(LSTM)[37] for sequence modeling, etc. This compartmentalized design has many flaws, making it difficult for the model to fully utilize global information and long-range dependencies. In recent years, the Transformer model has achieved excellent performance in sequence modeling tasks due to its powerful self-attention mechanism and end-to-end architecture design[4].It mainly uses an encoder-decoder architecture to achieve from image to text in OCR tasks.
@@ 3.1 Transformer-based OCR
The image encoder we adopt is based on the ViT model[6], which receives the input image $x_{\mathrm{img}}\in\mathbb{R}^{3\times H_{0}\times W_{0}}$ and resizes it to a fixed size $(H,W)$. Then, the input image is divided into $N$ image blocks of fixed size (P, P), tiled into vectors, and linearly projected to D-dimensional vectors. To preserve the position information of each pixel in the encoded image, absolute position embeddings are added to the embedding vector of each image patch. Next, the sequence of vectors is fed into a series of identical Transformer encoder layers. Each Transformer encoder layer contains a self-attention layer and a feed-forward layer. After each layer, a normalization layer and residual connections are applied. The self-attention mechanism calculates the similarity between each query and all keys, and uses the similarity as a weight to weight all values, and finally obtains the output. The feedforward layer usually consists of two fully connected layers and uses an activation function (such as ReLU and Swish[15]) for nonlinearization.
$\operatorname{Self-Att}(Q,K,V)$ $=V\operatorname{softmax}\big{(}\frac{K^{\top}Q}{\sqrt{D_{k}}}\big{)}$ (1)
$\operatorname{MultiHead}(Q,K,V)$ $=W_{o}[\operatorname{head}_{1};\cdots;\operatorname{head}_{h}]$ (2)
$\operatorname{head}_{i}$ $=\operatorname{Self-Att}(Q_{i},K_{i},V_{i})$ (3)
$\forall i\in\{1,\cdots,h\},\quad Q_{i}$ $=W_{q}^{i}Q,K_{i}=W_{k}^{i}K,V_{i}=W_{v}^{i}V$ (4)
@@
$y_{tw}$ $=Proj(Emb(w_{t}))$ (7)
$\text{Softmax}(y_{tw})$ $=\frac{e^{y_{tw}}}{\sum_{v=1}^{V}e^{y_{tv}}}\quad\text{for }w=1,2,...,V$ (8)
Where $V$ denotes the vocabulary size, $Emb(w_{t})$ denotes the embedding vector of the word $w_{t}$, and $Proj(\cdot)$ is the linear transformation that maps the embedding vector to the hidden state space.
@@ 3.2 Parameter-efficient Fine-tuning
$W=W_{0}+BA$ (9)
where $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are two small matrices, called LoRA weights, which are parameters that need to be learned during the fine-tuning process. $r$ is a hyperparameter that controls the rank of LoRA weight. When $r\ll\min(m,n)$, the number of parameters required by the LoRA weight will be much smaller than the original model weight matrix. By adding this low-rank update, LoRA enables efficient fine-tuning of pre-trained models. During the inference phase, the fine-tuned LoRA weights are combined with the original model weights to obtain the final inference weights. Since only two small matrices are stored, LoRA greatly reduces the additional parameter storage space required for fine-tuning. In our text decoder, we apply LoRA to both the query, key and value projection matrices of the self-attention module of the decoder and the encoder-decoder self-attention module, as well as the linear layer of the output. This parameter-efficient fine-tuning method allows us to significantly reduce the number of parameters required to fine-tune the text decoder while maintaining performance. Specifically, the low-rank matrix of LoRA is used to reconstruct the multi-head attention mechanism in the Transformer block, and its incremental update can be expressed as:
$\operatorname{head}_{i}=\operatorname{self-att}({Q}_{i},{K}_{i},{V}_{i})$ (10)
$Q_{i}=(W_{q,i}+B_{q,i}A_{q,i})\mathbf{x}$ (11)
$K_{i}=(W_{k,i}+B_{k,i}A_{k,i})\mathbf{x}$ (12)
$V_{i}=(W_{v,i}+B_{v,i}A_{v,i})\mathbf{x}.$ (13)
@@
$O=(W_{o}+B_{o}A_{o})\operatorname{Concat}(\operatorname{head}_{1};\cdots;\operatorname{head}_{h})$ (14)
@@
$\Delta W=\mathbf{u}\frac{V+\Delta V}{\|V+\Delta V\|_{c}}=\|W_{0}\|_{c}\frac{W_{0}+BA}{\|W_{0}+BA\|_{c}}$ (15)
@@
$Q_{i}$ $=(W_{q,i}+\|W_{q,i}\|_{c}\frac{W_{q,i}+B_{q,i}A_{q,i}}{\|W_{q,i}+B_{q,i}A_{q,i}\|_{c}})\mathbf{x}$ (16)
$K_{i}$ $=(W_{k,i}+\|W_{k,i}\|_{c}\frac{W_{k,i}+B_{k,i}A_{k,i}}{\|W_{k,i}+B_{k,i}A_{k,i}\|_{c}})\mathbf{x}$ (17)
$V_{i}$ $=(W_{v,i}+\|W_{v,i}\|_{c}\frac{W_{v,i}+B_{v,i}A_{v,i}}{\|W_{v,i}+B_{v,i}A_{v,i}\|_{c}})\mathbf{x}.$ (18)
@@
$\nabla_{A}\mathcal{L}=B^{\top}\frac{\partial\mathcal{L}}{\partial V^{\prime}},\nabla_{B}\mathcal{L}=\frac{\partial\mathcal{L}}{\partial V^{\prime}}A^{\top}$ (19)
$\frac{\partial{\mathcal{L}}}{\partial{V^{\prime}}}=\frac{\mathbf{u}}{\|V^{\prime}\|_{c}}\left(I-\frac{V^{\prime}V^{\prime\top}}{\|V\|_{c}^{2}}\right)\frac{\partial{\mathcal{L}}}{\partial{W}}$ (20)
@@
Table 1: Optimization of computational resources and model trainable parameters using LoRA, DoRA and DLoRA methods, which load the pre-trained weights of the model using half-precision (fp16).
@@
  • Full LoRA or DoRA: the encoder and decoder both adopt LoRA or DoRA method;
  • Hybrid configuration: the encoder uses DoRA method, while the decoder adopts LoRA method;
  • Full parameter fine-tuning: no PEFT method, all parameters participate in the training.
@@ 4.1 Experiment Setup
$\mathrm{Precision}$ $=\frac{\text{Correct matches}}{\text{The number of detected words}}$ (23)
$\mathrm{Recall}$ $=\frac{\text{Correct matches}}{\text{The number of ground truth words}}$ (24)
@@ 4.2 Experiment Result
Table 3: Different models are simultaneously evaluated in IAM,SROIE, STR benchmark and Mixed Dataset average CER and average F1 score.